OpenStudy (anonymous):

Let u = <-6, -2>, v = <-2, 3>. Find -3u + 2v.

OpenStudy (anonymous):

@phi @KingGeorge please I'm slowly dying

OpenStudy (kinggeorge):

So first, we just have to multiply the vectors by the scalars -3 and 2. First we have\[-3u=-3<-6,-2>=<-3(-6),-3(-2)>=<18,6>.\]Can you tell me what 2v is?

OpenStudy (anonymous):

I figured out the answer it is (14,12) right?

OpenStudy (kinggeorge):

That looks correct to me.

OpenStudy (anonymous):

YES thank you! Can you help me with one more please? (: @KingGeorge

OpenStudy (kinggeorge):

Sure.

OpenStudy (anonymous):

A tunnel is in the shape of a parabola. The maximum height is 16 m and it is 16 m wide at the base as shown below. What is the vertical clearance 7 m from the edge of the tunnel?

OpenStudy (kinggeorge):

Since it's a parabola, a quadratic equation will model that line very well. From the way it's drawn, we know the vertex is at (0,0), and it has a point at (16,-16). So can use the form\[y=a(x-h)^2+k\]where (h,k) is the vertex. So \(h=0\) and \(k=0\). Then we have that \[-16=a(16-0)^2+0=a16^2\implies a=-\frac{1}{16}\]So your quadratic equation will be \[-\frac{1}{16}x^2\]Make sense so far?

OpenStudy (anonymous):

yes!

OpenStudy (kinggeorge):

Now, when we're 7m from the edge, this corresponds to an x-value of 16-7=9. So plug in 9 for x to get\[-\frac{1}{16}9^2=-\frac{81}{16}\]But we want the distance of this from the base of the parabola, which is at -16. So we look at\[-\frac{81}{16}-(-16)=\frac{175}{16}=10.9375\]So this would be the vertical clearance they're looking for. Still make sense?